An annuity pays quarterly for 16 years at the beginning of each quarter. Payments during the first year are 250/quarter. Year 2 payments are each 105. The payments go up by 150 for each subsequent year but are level within each year. AEIR = 5%, please find the accumulated value of this annuity just prior to the payment at t=15.

A.

79,798.78

B.

72,848.29

C.

The answer does not appear here

D.

156,104.12

E.

154,014.68

and this is the previous question

An annuity pays quarterly for 16 years at the beginning of each quarter. Payments during the first year are 250/quarter. Year 2 payments are each 105. The payments go up by 150 for each subsequent year but are level within each year. AEIR = 5%, eventually we would like to find the accumulated value of this annuity just prior to the payment at t=15.

What is the quarterly interest rate i 1/4?

D.

.0122722355
Answer #1

The correct answer is the first option i.e. option A.   79,798.78

r = (1 + 5%)1/4 - 1 = 0.012272234

Please see the table below. Please be guided by the header to understand the mathematics. The last row highlighted in yellow is your answer. Figures in parenthesis, if any, mean negative values. All financials are in $.

t n = 60 - t Cash flows, C FV = C x (1 + r)^n
0 60 250          519.73
1 59 250          513.43
2 58 250          507.21
3 57 250          501.06
4 56 105          207.89
5 55 105          205.37
6 54 105          202.88
7 53 105          200.42
8 52 255          480.84
9 51 255          475.01
10 50 255          469.25
11 49 255          463.56
12 48 405          727.32
13 47 405          718.50
14 46 405          709.79
15 45 405          701.19
16 44 555          949.24
17 43 555          937.73
18 42 555          926.36
19 41 555          915.13
20 40 705      1,148.37
21 39 705      1,134.45
22 38 705      1,120.70
23 37 705      1,107.11
24 36 855      1,326.39
25 35 855      1,310.31
26 34 855      1,294.42
27 33 855      1,278.73
28 32 1005      1,484.84
29 31 1005      1,466.84
30 30 1005      1,449.06
31 29 1005      1,431.49
32 28 1155      1,625.20
33 27 1155      1,605.50
34 26 1155      1,586.03
35 25 1155      1,566.81
36 24 1305      1,748.82
37 23 1305      1,727.62
38 22 1305      1,706.68
39 21 1305      1,685.99
40 20 1455      1,856.99
41 19 1455      1,834.48
42 18 1455      1,812.24
43 17 1455      1,790.27
44 16 1605      1,950.89
45 15 1605      1,927.24
46 14 1605      1,903.87
47 13 1605      1,880.79
48 12 1755      2,031.63
49 11 1755      2,007.00
50 10 1755      1,982.67
51 9 1755      1,958.63
52 8 1905      2,100.26
53 7 1905      2,074.80
54 6 1905      2,049.65
55 5 1905      2,024.80
56 4 2055      2,157.75
57 3 2055      2,131.59
58 2 2055      2,105.75
59 1 2055      2,080.22
Total    79,798.78